OpenStudy (anonymous):

the measure of one angle is 24 degrees greater than the measure of its complement. What are the measures of the angles?

OpenStudy (anonymous):

let the angles be x and x+24 x+(x+24)=90 2x+24=90 2x=90-24 x=66/2 x=33 therefore,angles are 33 and 33+24=57

OpenStudy (anonymous):

x=(90-x) +24 ===>2x=114==>x=57 ==> the compliment=33

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

np :)